OpenStudy (anonymous):

You have to Divide. . s^2 -5s OVER s^2-s-20 / s-3 OVER s+4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

((s^2 -5s)/(s^2-s-20)) /((s-3)/(s+4)) =((s^2 -5s)/(s^2-s-20))*((s+4)/(s-3)) =[s*(s-5)/((s-5)(s+4)}*((s+4)/(s-3)) =s/(s-3)

OpenStudy (anonymous):

} s/b ]